OpenStudy (anonymous):

which expression is the simplified form of sqrt 108x^3 y^8 A. sqrt 5xy^4 sqrt3x B.6xy^4 sqrt3x C. 5x^2 y^8 sqrt3x D. 6x^2 y^8 sqrt3x

OpenStudy (anonymous):

I think it's B

OpenStudy (anonymous):

108 = 3^2 * 2^2 * 2 x^3 = x^2 * x y^8 = y^(2*4) So we have a 6xy^4 outside the square root and 3x inside the square root.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you're welcome